Solve this problem by using a two order system.
John is 5 years older than Mary. In 10 years, twice John's age decreased by Mary's age is 35, and John's age will be twice Mary's current age. Find their ages now.
John is years old ?, and Mary is ? years old.

Answer by richwmiller(17219)

Again no idea what you mean by "two order system"
j=m+5 and
2(j+10)-(m+10)=35 and
j+10=2m
j=20
m=15

We can solve by just using the first and last equations.
j=m+5
j+10=2m
m+5+10=2m
m+15=2m
15=m
j=20
Just for fun lets check using the unused equation
2(j+10)-(m+10)=35
2(30)-25=35
60-25=35
35=35
ok